Circular No. 8582

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF (87) SYLVIA F. Marchis, Department of Astronomy, University of California, Berke-

ley; and P. Descamps, D. Hestroffer, and J. Berthier, Institut de Mécanique Céleste et de Calcul des Éphémérides, Observatoire de Paris, report the discovery of a new satellite of (87) Sylvia (Marchis et al. 2005, Nature 436, 822). S/2004 (87) 1, of diameter perhaps 7 km, with maximum separation 0′′.44 from the primary, has a prograde near-circular orbit with estimated a = 710 km, P = 1.379 days. Another satellite, S/2001 (87) 1, of diameter perhaps 18 km, previously reported by M. E. Brown and J.-L. Margot (IAUC 7588), is estimated by Marchis et al. to have a = 1360 km, P = 3.650 days. Bearing in mind that the discoverer (Pogson 1866, MNRAS 26,

311) named (87) Sylvia for Rhea Sylvia (or Silvia), the mother of Romulus, with his twin brother the alleged founder of Rome, the IAU Committee on Small-Body Nomenclature has approved the permanent designations and names (87) Sylvia I = Romulus and (87) Sylvia II = Remus for the outer satellite S/2001 (87) 1 and the inner satellite S/2004 (87) 1, respectively.

NOVA IN THE SMALL MAGELLANIC CLOUD 2005 W. Liller, Viña del Mar, Chile, reports his discovery, on two Tech Pan

films taken on Aug. 6.388 UT with a 0.2-m Schmidt camera, of a possible nova (mag ≈ 10.4) located at α = 1^{h}15^{m}.0, δ = −73^{o}26′ (equinox 2000.0); a photo on Mar. 23 showed nothing at this position to mag 13. L. A. G. Monard, Pretoria, S. Africa, reports the following CCD position of the new object: α = 1^{h}14^{m}59^{s}.92, δ = −73^{o}25′35′′.8 (uncertainty ± 0′′.2; with magnitudes B = 12.0, V = 11.6, and R = 10.8 around Aug. 7.84).

E. Mason, European Southern Observatory (ESO); A. Brandeker, Department of Astronomy and Astrophysics, University of Toronto; A. Ederoclite, ESO; and M. Della Valle, Istituto Nazionale di Astrofisica, Osservatorio di Arcetri, confirm that the new object discovered by Liller is a classical nova in outburst. An échelle spectrogram (range 420–840 nm) of the nova was taken on Aug. 8.17 UT with the ESO New Technology Telescope (+ EMMI). A preliminary analysis shows that the spectrum is dominated by Balmer, Fe II (42 and 74), and Na I (1) emission lines. The Hβ emission line appears as strong as the Hα line. The line profile of the Balmer emissions is very broad and slightly asymmetric with a stronger blue peak. The average FWHM measured on the Balmer lines is ∼ 3200 km/s. There is a marginal evidence of a P-Cyg absorption on the blue wing of the Hβ emission, which possibly indicates an expansion velocity of ∼ 2100 km/s.

2005 August 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green